# A Causally Informed Pretraining Approach for Multimodal Foundation Models: Applications in Remote Sensing

### v3

@@ 8. Results: Pixel Wise Crop Mapping
Dataset and Problem Setting: Pixel wise crop mapping has been a topic of great study in the remote sensing community with numerous applications. Our data for finetuning comes from Sentinel2 and ERA5 land data for the region of the T11SKA Sentinel tile in the California Central Valley, a region rich in various crop classes and has been used for crop type mapping in various other works(Ghosh et al., 2021b, a; Ravirathinam et al., 2024) from the years 2018 and 2019. Like other works, we get our labels for this region from the Cropland Data Layer(CDL), an annually released land cover map for the entire continuous US by the USDA. A diagram of the CDL labels and geographic location of T11SKA tile can be seen in Appendix Figure 8.
In this problem setting, we use 10 spectral images (biweekly from May to Sept) and weather data (within that timeframe) to predict pixel wise crop labels at a 10m resolution. Similar to WSTATT (Ravirathinam et al., 2024), we adopt a grid based training method, by splitting the entire region into train, validation and test grids and also follow their preprocessing steps including combination and erosion. Note that the number of timestamps passed in the downstream task is different from the number passed during pretraining highlighting the temporal flexibility of our architecture. Due to the prediction related nature of this task, models from all 4 pretraining tasks are finetuned. A decoding head for semantic segmentation is appended and finetuned using cross entropy loss. Encoder is kept fixed. For more architecture and implementation details please refer to Appendix Section A.5.
2019 Test 11 class average F1 Scores
SM-MR MM-MR SM-VSF CI-VSF
Average 0.5331 0.5789 0.5731 0.6233
Performance Table 3 compares the performance of CI-VSF and other pretraining approaches when finetuned for crop mapping task. Though finetuning is done using 2018 data, testing is done on 2019 data, assessing the robustness of the approaches. We can see that CI-VSF performs better than other variants (Classwise scores shown in Appendix table 6). We also observe that MM-MR performs better than SM-MR, which is because inclusion of weather in input is in general better for crop mapping task (Ravirathinam et al., 2024), however MM-MR lacks capturing the relationship between modalities, thus its inferior performance when compared to CI-VSF. Finally, the ability of CI-VSF’s embeddings to generalize across years after finetuning indicates that the embeddings have captured crucial information that go beyond what is just present in spectral imagery.
@@ 9.1. Missing Image Prediction
Dataset and Problem Setting: With the vast amount of satellite imagery being captured on a daily basis, corrupted or missing data is an occurring phenomenon. In some cases, the areas of missing/corrupted data is provided as a mask, but in other cases this mask is not provided, making it hard to clean or filter data, thus leaving researchers to use these corrupted images in their work(Chen et al., 2022; Zhao et al., 2021). In this downstream task, we aim to fill these corrupted/missing values in the absence of these masks, i.e the methods have no access/information as to which pixels are corrupted, making it a very relevant downstream task.
@@
Performance: Table 4 compares the Mean Squared Errors across the various finetuned models. We can see that as the percentage of missing values increases, the MSE values go up, which can be expected due to the task getting harder. However, we can note that the CI-VSF pretrained models have significantly lower MSE values when compared to MR variants. We also note that CI-VSF performs the best across all missing value percentages, with the error not rising as much as the other variants. This is because the causal nature of CI-VSF helps in filling in the missing values. We also observe that the difference between SM-MR and MM-MR is not very high, thus furthering our hypothesis that MM-MR method of pretraining does not effectively capture the relationship between the different modalities of data.
@@
Table 4. Comparison on missing image prediction donwstream task across models finetuned from different pretraining tasks. Mean Squared Errors on various levels of missing image percentages shown
@@
% Missing SM-MR MM-MR SM-VSF CI-VSF
50% 792.68 788.94 362.02 326.43
70% 820.46 814.75 394.32 337.79
90% 826.23 820.43 404.32 343.88
In Figure 6 we compare some output images from our 50 percent missing values experiment across the four methodologies. Each row corresponds to a test sample, and a common trend across rows is that finetuned models pretrained using MR based methodologies do not fill the missing portions very well. In the first row, we can see that SM-VSF has added some false greenness but CI-VSF has not, showing that due to weather information, CI-VSF knows that growth has not occurred. Row 2 depicts a dried up river bed, but one can observe that SM-VSF fills the image with a wet river bed, whereas CI-VSF correctly estimates the dried up bed, once again proving that weather information has helped here. In the final row, we see a case where SM-VSF did not green up the region, but CI-VSF did, which once again is a result of causal embeddings.
9.2. Future Image Forecasting
Figure 7. Image Forecast Downstream Task comparison. Row 1 depicts a crop field, Green arrows depict regions of growth, and Red arrows depict regions of harvest, CI-VSF captures both these phenomena better than SM-VSF. Row 2 depicts a case where CI-VSF adds snowfall accurately compared to SM-VSF. Row 3 depicts a case where CI-VSF does not change land cover due to terrain but SM-VSF adds false greenness. Please zoom for better viewing. Larger version of this image is available in Appendix Figure 12
Dataset and Problem Setting: In this downstream task, we would aim to estimate a satellite image in the future given past data. Some applications where this task setting is applicable includes estimating how crops will grow in future (use the forecasted growth to calculate crop yield), predicting unavailable image due to revisit interval of satellite, etc. This downstream task is actually very similar to our pretraining objective, except here we focus only on the prediction of the last image of the series and do not include any masking of satellite imagery. So we can use the same training, validation and testing regions used in pretraining. However instead of using all images for finetuning, we randomly sample 1000 for training and 500 for testing and validation each. Like the soil moisture forecasting task, due to the forecasting related nature of this task, it would be inappropriate to finetune with the SM-MR and MM-MR embeddings. Decoder is reinitialised and finetuned using Mean Square Error loss. Encoder is kept fixed. For more architecture and implementation details please refer to Appendix Section A.7.
Performance:
@@ 9.2. Future Image Forecasting
Image Forecasting Downstream Task
Forecast Day Range SM-VSF CI-VSF
0 - 25 days 340.13 237.21
25 - 50 days 591.54 278.83
50 - 100 days 1093.62 358.23
More than 100 days 1112.84 457.27
@@
Table 5 compares the finetuned models performance on the image forecasting task, reporting average mean squared errors for different forecast ranges. We can observe that in the low forecast range (i.e 0 to 25 days into the future), CI-VSF fintuned model performs moderately better than the SM-VSF finetuned counterpart. However, as we increase the forecast day range, we can see the difference in performance between the two models diverge, with significant differences in beyond 100 day forecasts. This can be attributed to the fact that the encoder of CI-VSF has learnt the causal relationship between weather and satellite imagery and so is able to forecast well into the future with minimal finetuning. To highlight this further, some specific examples in the high forecast range are explored.
@@
Figure 7 shows a comparison of the forecasted images from these finetuned models,i.e SM-VSF and CI-VSF, on 3 independent examples. Each row correspond to a sample, with the first 6 images corresponding to the satellite component of the input series to the model, the weather component is not shown in the image but is passed along with the satellite component(as shown in Figure 4). Row 1 depicts an example of a crop field, with the final forecast image being 120 days following the 6th image in the input series. From the groundtruth image, we see that harvest has occurred in the circular fields and growth has happened in the top left corner field. Comparing the forecasted images from SM-VSF and CI-VSF, we can see that CI-VSF is able to capture both the harvest and the growth of the crops whereas SM-VSF is not able to capture these changes, showing that causal land cover dynamics have been captured by CI-VSF. Row 2 shows another example of a crop field, with the forecast image being 90 days in the future. We can see from the groundtruth image that 90 days later snow is present in the field, which is captured by CI-VSF but not SM-VSF, whose prediction shows a faded green field. This illustrates the ability of CI-VSF to capture the relationship between precipitation and temperature (i.e., precipitation during cold winter days can fall as snow). One can also notice that evergreen regions within the forecasted image of CI-VSF have less snowfall when compared to the fields, showing capture of terrain information. This is further reflected in Row 3, where a mountainous region is depicted and the forecast image 255 days in future. We can see that even after 255 days there is not much change in the region, which is correctly captured by our method (CI-VSF), illustrating terrain information capture, whereas SM-VSF seems to add some false greenness. We further observe that all images from the forecast of SM-VSF appear blocky, inspite of long periods of training. This is because the model is unable to forecast images without weather information.

### v4

@@ 7. Results: Pixel Wise Crop Mapping
Dataset and Problem Setting: Pixel wise crop type mapping has been a topic of great study in the remote sensing community with numerous applications. In this section, we explore the results of finetuning the models pretrained with the US Cropland Dataset for the task of crop type mapping. For results with the Global Dataset please refer to Appendix section A.4. Since we are dealing with the models from the US Cropland Dataset, we chose to finetune the model embeddings with regions in the 10 test tiles. In specific, we use just 400 regions of size 128x128 from the Sentinel Tiles of T11SKA, T16SBF, T15TUH, T13SGT in 2021 for finetuning the model’s embeddings for pixel wise crop type mapping(we chose these tiles due to their rich crop cover, and divided each tile into train, validation and test areas following a strategy similar to WSTATT(wstatt)). Like other works(statt; wstatt; calcrop21), we get our labels for these regions from the Cropland Data Layer(CDL), an annually released land cover map for the entire continuous US by the USDA. For each region we use 10 timestamps of data(approximately biweekly from the months of May to September) and the corresponding ERA5 weather data to create embeddings(Reasoning for choosing only 10 timestamps as opposed to using all year data is present in Appendix section A.4.3). For all the pretraining methods, we keep the layers used for creating the embeddings fixed, and add a decoding head with temporal attention for semantic segmentation. For all methods(KG-VSF, SM-MR, etc.) only the attention mechanism and decoder layers are finetuned using the 400 10-timestamps samples using cross entropy loss for 50 epochs each. Note that the number of timestamps passed in the downstream task is different from the number passed during pretraining highlighting the temporal flexibility of our architecture. For more architecture and implementation details please refer to Appendix Section A.4.
Table 1. Comparison on downstream task across 5 seeds of crop mapping across the pretraining tasks, Prithvi and crop type mapping models
Crop Type Mapping
Model Avg 10 class F1 Score
Model Mean(Std dev)
STATT 0.3032 (0.0249)
WSTATT 0.3289 (0.0240)
Prithvi EO.2 0.5050 (0.0189)
SMMR 0.6809 (0.0203)
MMMR 0.7056 (0.0169)
SMVSF 0.4407 (0.0224)
MMVSF 0.5426 (0.0122)
KGVSF 0.7602 (0.0139)
Performance Table 1 compares the performance of KG-VSF, other pretraining approaches, crop type mapping specific models(STATT(statt), WSTATT(wstatt)) and Prithvi EO.2(szwarcman2025prithvieo20versatilemultitemporalfoundation) when finetuned for crop mapping task using the 400 samples described above on 400 test samples(100 from each tile). In this table we compare the Average F1 Scores (along with standard deviations) across 10 crop classes for 5 seeds. One can observe that KGVSF performs the best across all methods and also shows relatively less variability across methods. We can observe that KGVSF outperforms both MMMR and MMVSF even though both methods also have weather as inputs for creating their embeddings, showing these pretraining methods lack capturing causal relationships between modalities leading to their inferior performance compared to KGVSF. One can also observe that STATT and WSTATT perform poorly, which is due to the low number of samples used for training the model(400 samples). We observed that if you increase the number of training samples used(around 4000), the performance of these models catches up to KGVSF(WSTATT reaches an F1 score of 0.84, MMMR reaches 0.89, and KGVSF reaches 0.92). From the table, we also see that Prithvi performs worse than SMMR. This could be due to a number of reasons including different resolutions of embeddings, and different pretraining datasets (Prithvi’s pretraining dataset may not be tuned for cropland areas unlike ours which focuses on crop lands). It is also shown that Prithvi just outperforms Unet, which STATT was shown to beat by a good margin. Overall, we can see that embeddings from KGVSF are able to be finetuned for crop type mapping using little samples and obtain good performance relative to other pretraining methods and models.
8. Results: Soil Moisture Tasks
In this section, we compare the performance of fine-tuned models on soil moisture related tasks, specifically soil moisture prediction and soil moisture forecasting.
8.1. Soil Moisture Estimation
Dataset and Problem Setting: Estimation of Soil Moisture at current timestamp using radar imagery has been a topic of interest in the remote sensing community(adab2020machine; lee2019estimation; senanayake2021estimating). Typically, multiple sources of data including satellite data, radar data, weather data, depth data, etc. are used to estimate in-situ measurements from their site where groundtruth readings are available. In our downstream task, we aim to emulate this task and demonstrate how a pretraining task that captures causal relationships in its embeddings is better for the task of soil moisture estimation. In this section, we explore the results of finetuning the models pretrained with the US Cropland Dataset for this task. For results with the Global Dataset please refer to Appendix section A.5. We focus on 6 Sentinel tiles (T11SKA, T10SEJ, T14SKC, T15TUH, T16SBF, T14RQT) regions (geographic locations shown in Appendix Figure 14) and obtain the corresponding Sentinel2, ERA5 and SMAP (’soil_moisture_am’ band) values for those regions at their optimal temporal and spatial resolutions. These SMAP values serve as our groundtruth for soil moisture(li2021improved), i.e our goal would be to use the Sentinel and ERA5 data to predict SMAP value (’soil_moisture_am’). The utility of this task would be in cases where SMAP values are not present or the SMAP satellite fails and we need to estimate the SMAP values for the missing days(this is possible since we would have satellite imagery and weather data for these days). In this section, we explore the results of finetuning the models pretrained with the US Cropland Dataset for this task, for results with the Global Dataset please refer to Appendix section A.5.
In our problem setting, the input would be a time series of either one or multiple modalities to get embeddings from the respective pretraining objectives. Then using these embeddings we would aim to estimate the soil moisture at that location for each timestamp. We use a timeframe of 6 satellite images here and aim to estimate the soil moisture at each timestamp. Each tile is split into 100 grids and a 60-20-20 train-val-test split is done, ensuring no overlap in regions. Multiple samples are generated from each grid, and after preprocessing, we retain 300 samples(50 from each tile) for training, 150 for validation(25 from each tile) , and 150 for testing(25 from each tile). Models from all 4 other pretraining tasks are finetuned by adding decoding head for prediction. Note that for all methods, layers used to create embeddings are fixed and only the decoder is finetuned using Mean Absolute Error loss. For more architecture and implementation details please refer to Appendix Section A.5.
Table 2. Comparison of models fine-tuned across 5 seeds from different pretraining tasks on the downstream task of soil moisture prediction. R Square Score for each experiment setting is shown.
Soil Moisture Prediction
R Square
Model Mean (Std dev)
Satellite only 0.1249 (0.0532)
Weather only 0.4813 (0.0101)
Weather + Satellite 0.4906 (0.0110)
Prithvi EO.2 0.3377 (0.0098)
SMMR 0.1845 (0.0393)
MMMR 0.6039 (0.0148)
SMVSF 0.2661 (0.0180)
MMVSF 0.7606 (0.0162)
KGVSF 0.8081 (0.0150)
Performance: The performance of finetuned models across 5 seeds on this task are presented in Table 2. We include comparison with Prithvi EO.2(szwarcman2025prithvieo20versatilemultitemporalfoundation) and other baseline implementations including a models trained from scratch that uses only weather, only satellite data and one model that uses both sources of data. We can see that KGVSF outperforms other methods, once again beating both MMMR and MMVSF (both methods also have weather as inputs for creating their embeddings), showing that performance gain of KGVSF is coming from causal capture within the embeddings. One can also observe methods that rely on only satellite imagery(Eg. Prithvi, SMMR, SMVSF) perform suboptimally, which can be explained as satellite imagery being less linked to soil moisture compared to weather. Overall, we can see the benefit of KGVSF embeddings in this task as well.
8.2. Soil Moisture Forecasting
Dataset and Problem Setting: Soil moisture forecasting plays an important role in the remote sensing domain, enabling various applications such as optimizing crop management, planning irrigation schedules, and improving drought monitoring and mitigation strategies (dubois2021short; togneri2022soil). In this downstream task, we aim to evaluate the ability of model’s embeddings to forecast soil moisture estimates. Specifically, we seek to predict soil moisture at a given location at an arbitrary future time based on embeddings from past satellite imagery and weather and weather data up to that future time. The embeddings come from each pretraining strategy. We focus on the same 6 Sentinel tiles (T11SKA, T10SEJ, T14SKC, T15TUH, T16SBF, T14RQT) regions (geographic locations shown in Appendix Figure 14) and obtain the corresponding Sentinel2, ERA5 and SMAP (’soil_moisture_am’ band) values for those regions at their optimal temporal and spatial resolutions. Our goal is to forecast SMAP (’soil_moisture_am’ band) values at some point in the future using embedding from past context and weather upto forecast date. In this section, we explore the results of finetuning the models pretrained with the US Cropland Dataset for this task. For results with the Global Dataset please refer to Appendix section A.6.
To estimate a soil moisture value in the future, we provide past satellite image time series and weather data to get embeddings from each pretraining objective. Then using those embeddings and weather till a future date we estimate the soil moisture at a future date. We use an input timeframe of 6 satellite images and predict the soil moisture value at a random 7th timestamp in the future. Each tile is split into 100 grids and a 60-20-20 train-val-test split is done, ensuring no overlap in regions. Multiple samples are generated from each grid, and after preprocessing, we retain 300 samples for training, 150 for validation, and 150 for testing. To enhance model generalization, we ensure a diverse distribution of forecast range (i.e. the number of days into the future for which forecasts are made), allowing the model to learn both short-term and long-term forecasting patterns. Since we are forecasting using embeddings from past data and future weather data, we can test this task on all strategies of pretraining. A decoding head for forecasting is appended and finetuned using Mean Absolute Error loss. Encoder is kept fixed. For more architecture and implementation details please refer to Appendix Section A.6. To reiterate, we use embeddings from each pretraining strategy for capturing past information and use that for forecasting, so the focus of this task is on how the embedding from each pretraining strategy affects in forecasting task.
Table 3. Comparison of models fine-tuned from different pretraining tasks on soil moisture forecasting downstream task. Table reports average R Squared Score of forecasted SMAP values for different forecast day ranges.
Soil Moisture Forecasting Downstream Task
(R Squared Scores Reported (higher the better))
Forecast Day Range Weather Satellite SMMR MMMR SMVSF MMVSF KGVSF
0-25 days 0.5047 0.3821 0.5429 0.4100 0.6639 0.7834
25-50 days 0.4992 0.3651 0.5341 0.3986 0.6142 0.7034
50-100 days 0.4942 0.3519 0.4982 0.3821 0.5534 0.6532
100+ days 0.4931 0.3409 0.4465 0.3619 0.5139 0.6164
Performance: We train a model for each of the pretraining task settings using all tiles and a diverse range of forecast day (i.e varying number of days into the future for forecasting). To compare the performance of these models systematically, we compare predictions made across different forecast ranges in the test grids, reported in Table 3. From the Table, we can see that across all forecast day ranges KG-VSF’s finetuned model performs the best. We can that in the initial short day ranges, the performance is quite seperated but as we go larger forecast day ranges, we can see that performance saturates. This is caused due to the low number of samples used for finetuning, as the forecast day range increases, the effect of the embedding woudl diminish and the model would rely on the future weather to predict the soil mositure at that point. Nevertheless, we see that KGVSF performs the best showing that an embedding that captures causal information is beneficial.
@@ 9.1. Missing Image Prediction
Dataset and Problem Setting: With the vast amount of satellite imagery being captured on a daily basis, corrupted or missing data is an occurring phenomenon. In some cases, the areas of missing/corrupted data is provided as a mask, but in other cases this mask is not provided, making it hard to clean or filter data, thus leaving researchers to use these corrupted images in their work(chen2022joint; zhao2021evaluation). In this downstream task, we aim to fill these corrupted/missing values in the absence of these masks, i.e the methods have no access/information as to which pixels are corrupted, making it a very challenging task. It therefore tests whether the pretrained embeddings can infer physically plausible content under uncertainty, a realistic use case for cloud-removal or data-gap filling where clouds obscure parts of the image or no cloud mask is available. This setting is thus more challenging and practically meaningful than the masked-reconstruction tasks where the masked portions (e.g., a cloud mask) are known. In this section, we explore the results of finetuning the models pretrained with the Global Dataset for this task.
@@
Performance: Table 4 compares the Mean Squared Errors across the various finetuned models. We can see that as the percentage of missing values increases, the MSE values go up, which can be expected due to the task getting harder. However, we can note that the KG-VSF pretrained models have significantly lower MSE values when compared to MR variants. We also note that KG-VSF performs the best across all missing value percentages, with the error not rising as much as the other variants. This is because the causal nature of KG-VSF helps in filling in the missing values. We also observe that the difference between SM-MR and MM-MR is not very high, thus furthering our hypothesis that MM-MR method of pretraining does not effectively capture the relationship between the different modalities of data.
@@
In Figure 5 we compare some output images from our 50 percent missing values experiment across the four methodologies. Each row corresponds to a test sample, and a common trend across rows is that finetuned models pretrained using MR based methodologies do not fill the missing portions very well. In the first row, we can see that SM-VSF has added some false greenness, and MM-VSF does capture greenness, but KG-VSF does not make these mistakes, showing that the embeddings have captured how weather information leads to growth. Row 2 depicts a dried up river bed, but one can observe that SM-VSF and MM-VSF fill the image with a wet river bed, whereas KG-VSF correctly estimates the dried up bed, once again proving embeddings have captured relevant causal satellite-weather information. In the final row, we see a case where SM-VSF and MM-VSF do not green up regions, but KG-VSF did, which once again is a result of causal embeddings.
Table 4. Comparison on missing image prediction downstream task across models finetuned from different pretraining tasks. Mean Squared Errors on various levels of missing image percentages shown
@@
% Missing SM-MR $M$-MR SM-VSF $M$-VSF KG-VSF
50% 792.68 788.94 362.02 347.21 326.43
70% 820.46 814.75 394.32 367.82 337.79
90% 826.23 820.43 404.32 378.13 343.88
9.2. Zero Shot Future Image Forecasting
Dataset and Problem Setting: In this downstream task, we would aim to estimate a satellite image in the future given past data. Some applications where this task setting is applicable includes estimating how crops will grow in future (use the forecasted growth to calculate crop yield), predicting unavailable image due to revisit interval of satellite(Sometimes satellites do not revisit for upto 15 days, leaving researchers without a way to obtain a good estimate of the imagery in that timeframe), etc. This task requires understanding how weather interacts with land cover or else the models would rely simply on correlation between available past imagery. The pretraining objective of KGVSF is designed to capture this interaction. This downstream task is actually very similar to our pretraining objective, except here we focus only on the prediction of the last image of the series and do not include any masking of satellite imagery with no finetuning, hence the name zero shot. In this section, we explore the results of the models pretrained with the Global Dataset for this task. Since there is no overlap, we can use the same training, validation and testing regions used in pretraining. Due to the zero shot forecasting related nature of this task, it would be inappropriate to use the SM-MR and MM-MR pretrained models. All pretrained frameworks are kept fixed for the prediction of future image. For more architecture and implementation details please refer to Appendix Section A.8.
Figure 6. Image Forecast Downstream Task comparison. Row 1 depicts a crop field, Green arrows depict regions of growth, and Red arrows depict regions of harvest, KG-VSF captures both these phenomena better than SM-VSF. Row 2 depicts a case where KG-VSF adds snowfall accurately compared to SM-VSF. Row 3 depicts a case where KG-VSF does not change land cover due to terrain but SM-VSF adds false greenness. Please zoom for better viewing.
@@ 9.2. Zero Shot Future Image Forecasting
Zero Shot Image Forecasting Downstream Task
Forecast Day Range SM-VSF $M$-VSF KG-VSF
0 - 25 days 340.13 310.45 237.21
25 - 50 days 591.54 565.21 278.83
50 - 100 days 1093.62 1004.21 358.23
More than 100 days 1112.84 1097.67 457.27
@@
Performance: Table 5 compares the finetuned models performance on the image forecasting task, reporting average mean squared errors for different forecast ranges. We can observe that in the low forecast range (i.e 0 to 25 days into the future), KG-VSF finetuned model performs moderately better than the SM-VSF finetuned counterpart. However, as we increase the forecast day range, we can see the difference in performance between the two models diverge, with significant differences in beyond 100 day forecasts. This can be attributed to the fact that the encoder of KG-VSF has learnt the causal relationship between weather and satellite imagery and so is able to forecast well into the future with minimal finetuning. To highlight this further, some specific examples in the high forecast range are explored.
@@
Figure 6 shows a comparison of the forecasted images from these models,i.e SM-VSF, MM-VSF and KG-VSF, on 3 independent examples. Each row correspond to a sample, with the first 6 images corresponding to the satellite component of the input series to the model, the weather component is not shown in the image but is passed along with the satellite component(as shown in Figure 4) for both MM-VSF and KG-VSF. Row 1 depicts an example of a crop field, with the final forecast image being 120 days following the 6th image in the input series. From the groundtruth image, we see that harvest has occurred in the circular fields and growth has happened in the top left corner field. Comparing the forecasted images from SM-VSF, MM-VSF and KG-VSF, we can see that KG-VSF is able to capture both the harvest and the growth of the crops whereas SM-VSF and MM-VSF are not able to capture these changes, showing that causal land cover dynamics have been captured in the embeddings of KG-VSF. Row 2 shows another example of a crop field, with the forecast image being 90 days in the future. We can see from the groundtruth image that 90 days later snow is present in the field, which is captured by KG-VSF but not the others, whose predictions show a faded green field. This illustrates the ability of KG-VSF to capture the relationship between precipitation and temperature on land cover (i.e., precipitation during cold winter days can fall as snow). One can also notice that evergreen regions within the forecasted image of KG-VSF have less snowfall when compared to the fields, showing capture of terrain information. This is further reflected in Row 3, where a mountainous region is depicted and the forecast image 255 days in future. We can see that even after 255 days there is not much change in the region, which is correctly captured by our method (KG-VSF), illustrating terrain information capture, whereas SM-VSF seems to add some false greenness. We further observe that images from the forecast of SM-VSF and MM-VSF might appear blocky, inspite of long periods of training. This is because the models are unable to forecast images into the future accurately without weather information.
10. Analysis of Embeddings
To assess the quality of each pretrained model we visualize year-end embeddings produced by each pretrained model (SM-MR, MM-MR, SM-VSF, MM-VSF, and KG-VSF) for selected crop types from the US Cropland Dataset for which USDA label quality is known to be generally good. First we resample USDA crop data layer (CDL) labels for 30-meter Landsat pixels to 10-meter Sentinel pixels. Recall that each 128×128 region is divided into 256 patches each of size 8x8 (section 6.2). We then retain only those patches that have at least 80% purity with respect to a single CDL crop class (i.e., one CDL crop label occurs in at least 80% of the 64 pixels). The embedding produced for each of these retained 8x8 patches by each pretrained model is projected into two dimensions using t-SNE, and each point is colored according to that patch’s CDL land-cover label. This allows us to directly visualize and compare how well each pretraining approach organizes different land-cover types in its learned latent space in Figure 7.
Figure 7. Comparison of embedding quality from the various pretraining objective via t-SNE plots. Embeddings from specific land cover classes are projected into 2d space via tsne. One can observe that KGVSF forms more seperable clusters relative to other methods.
We observe that KG-VSF forms tight, semantically consistent clusters with clear separation between crop types. In contrast, MR-based models produce more diffuse and overlapping clusters, particularly among phenologically similar crops, while VSF variants without knowledge guidance show weaker boundaries.
Overall, these t-SNE visualizations reveal that KG-VSF learns more structured and interpretable representations, organizing patches according to both land-cover semantics and temporal dynamics. This structure emerges purely from the pretrained embeddings, confirming that knowledge-guided pretraining captures meaningful cross-modal relationships even before task-specific fine-tuning.
11. Does KGVSF actually Capture Causality?
To verify that KGVSF is actually capturing causality and not just sharpening correlations, we performed two experiments. In the first experiment, keeping image history fixed we feed counterfactual weather and observe if the forecasted image changes in the expected direction. In the second experiment, we mute and shuffle weather and observe the performance on downstream tasks of crop type mapping and soil moisture prediction.
11.1. Counterfactual Weather Probing Test
To verify that KGVSF is capturing causality and not just sharpening correlation, in this experiment, we hold the input image history fixed and feed several counter-factual weather sequences and see if the predicted forecast image changes in the expected direction. We chose three samples from the US crop land areas and fed these samples with dry hot weather from Texas (southern latitude), cold wet (snowy) weathers from Northern MN (Northern latitude), and sunny temperate weather from California(middle latitudes) and visualised the outputs produced by model pretrained by KGVSF in Figure 8.
Figure 8. Predicted future images under different weather conditions for locations in Tennessee (Row 1), MN (row 2 and 3)
From the figure we can see that under different weather conditions, the forecasted image changes in the physically expected direction. One can notice that under dry hot weather conditions, all forecasted imagery lack greenery. Under the sunny temperate weather condition, forecasted images show greenery. Under the cold wet snowy weather, snow is added to all the images. Note that snow is not added throughout the image but rather added in the places where one would expect it. For example, in row 1, snow is added in the open farm land area but not on the parts (left portion of the location) that are covered by trees. Also snow is not added in the green region surrounding the river that may have a slope, making it harder for snow to accumulate there.
11.2. Mute and Shuffle Weather Test
We observed from the various downstream tasks that embeddings from KGVSF show better performance than embeddings from other pretraining schemes. To investigate if KG-VSF is benefitting from the causal signal itself or simply from having a second modality at the decoder, we conducted an additional ablation experiments on the pixel-wise crop mapping and soil moisture prediction tasks under two settings: (1) weather is muted (ie. all values are made zero) and (2) weather is shuffled as described below:
  • Zeroed-Out Weather: In this variant, we replace all weather values with zeros, effectively removing the contribution of the weather modality while keeping its structure in the input pipeline. This allows us to verify whether the improvement in the full model arises from the mere presence of a second modality in the decoder. Since the zeroed-out weather input provides no information, any improvement in performance beyond this variant can be attributed to the actual semantic value of the weather data rather than its structural inclusion.
  • Shuffled Weather: In this setting, we randomly shuffle the weather series across the year for each spatial location. This destroys the temporal and semantic alignment between the weather and the corresponding satellite imagery. As a result, the weather values lose their physical meaning with respect to the visual dynamics in the imagery, preventing the model from leveraging any true causal linkage between the two modalities. Consequently, we expect the embeddings to become suboptimal, leading to degraded performance when fine-tuned for downstream tasks.
Table 6. Comparing the performance of shuffling and muting weather experiments on the tasks of crop type mapping and soil moisture estimation.
Crop Type Mapping Soil Moisture Prediction
Variant Avg F1 Score R Square
Zero out weather 0.6399 0.3143
Shuffle weather 0.6933 0.3751
Normal 0.7602 0.8081
Table 6 compares the performance of these two experiments along with the performance of the base case with correct weather for both pixel wise crop mapping and soil moisture estimation. Across both tasks, we consistently observe that the performance of the shuffled and zeroed-out variants drops substantially compared to the normal case. The shuffled-weather case performs better than the zeroed-out one, likely because it still contains some information about the actual weather (e.g., average temperature, total amount of precipitation). Note that the drop is much steeper for the soil moisture task than for the crop type mapping task. We believe that this is due to the fact that the KGVSF embeddings are capturing several distinct signals: 1. Patterns in the satellite images, 2 patterns in the weather, and 3 causality between weather conditions and satellite images. For the crop mapping task, the first component alone likely has a lot of relevant information (as shown in our earlier results on SMMR) although components 2 and 3 also have additional values. However the soil moisture depends heavily on the weather (specifically precipitation), and thus muting or shuffling it causes huge degradation in the performance.
Overall these results demonstrate that KG-VSF’s performance gain does not arise from simply having an additional modality at the decoder. Instead, the improvement is due to the model’s ability to leverage physically meaningful, causally aligned weather information during pretraining, resulting in embeddings that encode the true driver-response relationships between weather and surface dynamics.
12. Evaluating Scalability of KGVSF
To assess the scalability of our approach and estimate potential downstream gains with larger-scale pretraining, we used the US Cropland dataset spanning 50 Sentinel-2 tiles across the United States as described in Section 5.2. As mentioned in Section 5.2, these tiles were divided into 30 for training, 10 for validation, and 10 for testing, and from each tile, approximately 3,000 spatiotemporal samples were generated using the 7-timestamp Sentinel-2 input configuration.
To systematically explore the impact of scale, we created progressively larger subsets of the training data by sampling an equal number of examples from each of the 30 training tiles: 500 samples per tile (15,000 total), 1,000 per tile (30,000 total), 1,500 per tile (45,000 total), and 2,000 per tile (60,000 total). For each dataset, we trained a model using the KG-VSF pretraining task under identical hyperparameter settings. The resulting pretrained embeddings were then fine-tuned on two representative downstream tasks: pixel-wise crop mapping and soil-moisture prediction.
As shown in Table 7, both downstream metrics improve consistently as the number of pretraining samples increases. This steady upward trend indicates that additional unlabeled Sentinel-2 data would likely yield further performance gains, reinforcing the scalability of the KG-VSF pretraining paradigm. Although computational limitations restricted our experiments to 60,000 samples, the results suggest further improvements are possible with larger datasets that are available publicly from Sentinel-2.
Table 7. Performance of KGVSF model pretrained with different number of samples on crop type mapping and soil moisture prediction.
Scalability Test Crop Type Mapping Soil Moisture Prediction
No of samples used in Pretraining KGVSF Avg F1 score R Square
15000 0.6496 0.6684
30000 0.7336 0.7308
45000 0.7521 0.7889
60000 0.7602 0.8081